OpenStudy (anonymous):

If the scale factor of two similar solids is 4:17, what is the ratio of their corresponding areas and volumes? A. 4:289 and 17: 4,913 B. 8:34 and 12:51 C. 16:289 and 64:4,913 D. 64:4,913 and 16:289

OpenStudy (anonymous):

take a cube 4*4*4 area =4*4=16 vol =4*4*4=64 " " " 17*17*17 area =17*17=289 vol =17*17*17=4913 answer C area(16:289) vol (64:4913)